# LLM4WM: Adapting LLM for Wireless Multi-Tasking

### v1

@@ IV-E Training Configuration
where $f_{loss,n}$ denotes the loss function for task $n$. And they are linearly combined with task weightings $\omega_{n}$. To ensure that all tasks are well trained, we use the Dynamic Weight Average (DWA) algorithm [32] to dynamically adjust each task’s weight based on its loss every epoch. The selection of loss function $f_{loss,n}$ fully considers the characteristics of the task itself. For classification problems, such as ${BF}$, we use the cross entropy loss function, and for regression problems, such as ${CP}$, we use normalized mean square error (NMSE) [22] as the loss function.
@@ V Experiments
In this section, we outline the simulation settings and subsequently evaluate the performance of the proposed LLM4WM method from several perspectives: overall performance, generalization ability, and stability. Comprehensive ablation experiments are also conducted to demonstrate the effectiveness of each module within the framework.
@@ V-A1 Datasets
We adopt the widely used channel generator QuaDRiGa [33] to simulate time-varying CSI datasets compliant with 3GPP standards. We consider a dual-frequency wireless system having a $1.9$ GHz sub-6G link and a $28$ GHz mmWave link. LOS propagation environment is assumed between BS and UEs such that the optimal transmitting beam for the mmWave link is determined by LOS direction. The hyper-parameters of the dataset generation are presented in Tab. II.
@@
In contrast, the mmWave link employs the TDD mode. For the sub-6G assisted mmWave beamforming task, the downlink analog precoding is also derived based on the spatial correlation of the uplink sub-6G channel estimated by the uplink pilot. The initial position of the user is randomized and the motion trajectory is set as linear type at a speed of $30$ km/h. The dataset consists of $20000$ samples, which are partitioned into a training set, validation set, and test set in a ratio of $9:1:2$. Specifically, the training set contains $15000$ samples, the validation set comprises $1600$ samples, and the test set includes $3400$ samples.
@@ V-A2 Baselines
  • FIFS [35]: FIFS is a CSI-based fingerprinting system that presents a coherence bandwidth-enhanced probability algorithm with a correlation filter to map objects to the fingerprints. We implement it for the radio environment mining tasks.
@@
  • MLP [36, 37]: Multi-Layer Perceptron(MLP) is a type of feedforward neural network commonly used for classification and regression tasks. Many works have employed MLPs to model complex mapping relationships in communication problems. We implement an MLP for the task of radio environment sensing and beam management.
@@
  • CNN [23]: In [23], a CNN-based predictor is proposed for FDD systems, treating the prediction process of time-frequency CSI data as a two-dimensional image processing task. It contains ten convolutional layers, where the convolution kernel size is $3\times 3$. We implement it for processing channel reconstruction tasks.
@@
  • Transformer [22]: In [22], a transformer-based parallel channel predictor is proposed for TDD systems to avoid error propagation problems. We implement it using $3$ encoders and $2$ decoders for processing channel reconstruction tasks.
@@ V-A3 Network and Training Parameters
In the simulation, we adopt the Multi-Task Adapter module with $N_{a}=8$ for input and output feature alignment. For the configuration of the MoE-LoRA fine-tuning method, we choose the number of experts to be $8$, as the number of experts is generally slightly greater than the number of tasks. Additionally, we set $r=8$ for each LoRA matrix. For the output module, as mentioned above, for the specific task, we use either a three-layer MLP with 768-dimensional features or a three-layer CNN with $33x$ 3 kernels for the feature process. and we employ only one single-layer fully connected network to align the output dimensions. The smallest version [42] of GPT-2 with $F=768$ feature dimension is adopted, the first $N_{L}=6$ layers of which are deployed. Both the warm-up and cosine annealing scheduler are employed to train LLM4WM. The first $50$ epochs serve as the warm-up phase, where the learning rate increases linearly from the minimum value of $1\times{10}^{-5}$ to $1\times{10}^{-3}$. In the subsequent training phases, the learning rate is dynamically adjusted using the cosine annealing scheduler. Several other hyper-parameters for model training are illustrated in Tab. III.
@@ V-A4 Performance Metric
In order to comprehensively evaluate the performance of the proposed scheme, we adopt an extra metrics, namely SE which reflect the performance of the communication system. SE is an important metric that reveals the achievable rate of the system, reflecting the effectiveness of communication. It is calculated by Eq. (4), where $\bm{h}_{t,k}$ is the actual CSI and $\bm{w}_{t}$ is obtained as Eq. (7) with predicted $\bm{h}_{t,k}$. The communication SNR is defined as $1/{\sigma_{n}^{2}}$ and set as $10$ dB.
@@ V-B1 Overall Performance
TABLE IV: Performance of LLM4WM and other baselines: For mmWave link, the maximum SE is $9.32$ $\cdot$$(\text{s}\cdot\text{Hz})^{-1}$; For sub-6G link, the maximum SE is $6.33$ $\cdot$$(\text{s}\cdot\text{Hz})^{-1}$. The best results are highlighted in bold, while the second-best results are underlined.
@@ V-B2 Generalization Experiments
TABLE V: Generalization Performance of LLM4WM and Other Baselines: the average loss across tasks is computed and presented in the final column labeled ”AVG.”. The best results are highlighted in bold, while the second-best results are underlined.
@@
Generalization is the model’s ability to maintain performance in new communication scenarios, crucial for real-world deployment as it reduces the need for frequent updates. We use only $10\%$ of the RMa dataset to transfer a model trained in the UMa scenario, as well as a model trained on the $1.9$ GHz sub-6 GHz link dataset to the $2.4$ GHz sub-6 GHz link dataset. Results in Tab. V indicate that, despite the challenges of multi-task generalization and transfer, our approach consistently outperforms others across most tasks. The slight dip in radio environment mining performance is due to the task’s relative simplicity in the LOS scenario, where smaller models like WiT excel. However, our model excels in more complex tasks like channel estimation, which requires understanding multidimensional features, further confirming that large models are better suited for dynamic real-world communication scenarios.
@@ V-B3 Hyper-parameter Analysis
To illustrate the rationale behind the hyperparameter settings, we conduct a thorough experiment into the impact of hyperparameters on the performance of LLM4WM. Specifically, we examine the effects of varying the LoRA rank and the number of experts, as depicted in Fig. 7. When we increase the LoRA rank while keeping the number of experts fixed at $8$, the performance of LLM4WM gradually improves. This can be attributed to the enhanced adaptability of the model to the data distribution due to the increase in trainable parameters. However, this improvement comes at the cost of increased training overhead. After weighing the trade-off between performance and computational cost, we determine that a LoRA rank of $8$ provided the optimal balance. Subsequently, with the LoRA rank fixed at its optimal value of $8$, we incrementally increase the number of experts. A similar trend to that of the LoRA rank was observed, as an increase in the number of experts effectively enhanced the model’s analytical and representational capacity. Balancing performance and computational efficiency, we determine that setting the number of experts to $8$ was the most appropriate choice.
@@ V-B4 Ablation Experiments
To validate the proposed modules, we conducted ablation experiments by removing or altering configurations of the multi-task adapter and backbone LLM modules. Variations for the multi-task adapter include w/o $\text{Adapter}_{\text{in}}$ (placing adapter only on the output side of the LLM), w/o $\text{Adapter}_{\text{out}}$ (placing adapter only on the input side of the LLM), and w/o Adapter (no adapters). For the backbone LLM, variations include w/o LLM (removing the large model) and frozen LLM (freezing pre-trained weights). Results in Table V show that all ablation configurations led to performance declines, highlighting the effectiveness of both the multi-task adapter and backbone LLM modules. Notably, the impact of removing the backbone LLM is significantly greater, indicating its critical role in the success of multi-task joint learning for wireless tasks.
@@ V-B5 Efficiency Evaluation
We compare the model training and inference cost of the proposed method with other baselines to assess the difficulty of deploying the model in practical scenarios, as shown in Tab. VII. All experiments are conducted on the same machine with $4$ Intel Xeon Platinum 8375C CPUs, $4$ NVIDIA GeForce RTX4090 GPUs, and $256$ GB of RAM. To facilitate representation and comparison, the data in the table reflects the average performance across tasks for each method. Notably, the MoE-LoRA fine-tuning method results in LLM4WM having trainable parameters comparable to those of smaller models. This highlights both the training efficiency and parameter efficiency of LLM4WM, as adding new tasks increases the model’s parameters by only about $1.13$ M, a small fraction of the total. Additionally, utilizing a lightweight backbone LLM ensures that the overall inference speed of LLM4WM remains acceptable. Thus, LLM4WM demonstrates significant potential for deployment in future communication scenarios marked by increasing demand and the need for customized services involving numerous tasks.

### v2

@@ IV-E Training Configuration
where $f_{loss,n}$ denotes the loss function for task $n$. And they are linearly combined with task weightings $\omega_{n}$. To ensure that all tasks are well trained, we use the Dynamic Weight Average (DWA) algorithm [33] to dynamically adjust each task’s weight based on its loss every epoch. The selection of loss function $f_{loss,n}$ fully considers the characteristics of the task itself. For classification problems, such as ${BF}$, we employ the cross-entropy loss function, while for regression problems, such as ${CP}$, the normalized mean square error (NMSE) [23] is used as the loss function.
@@ V Experiments
In this section, we first describe the simulation settings and then assess the performance of the proposed LLM4WM method from multiple perspectives: overall performance, generalization ability, and stability. We also conduct comprehensive ablation experiments to highlight the contribution of each module within the framework.
@@ V-A1 Datasets
We adopt the widely used channel generator QuaDRiGa [34] to simulate time-varying CSI datasets compliant with 3GPP standards. We consider a dual-frequency wireless system having a $1.9$ GHz sub-6G link and a $28$ GHz mmWave link. The hyper-parameters of the dataset generation are presented in Tab. II.
@@
In contrast, the mmWave link employs the TDD mode. For the sub-6G assisted mmWave beamforming task, the downlink analog precoding is also derived based on the spatial correlation of the uplink sub-6G channel estimated by the uplink pilot. The initial position of the user is randomized and the motion trajectory is set as linear type at a speed of $30$ km/h. The dataset contains a total of $20000$ samples. Specifically, the training set includes $15000$ samples, the validation set has $1600$ samples, and the test set consists of $3400$ samples.
@@ V-A2 Baselines
  • FIFS [36]: FIFS is a CSI-based fingerprinting system that introduces a coherence bandwidth-enhanced probability algorithm, utilizing a correlation filter to map objects to fingerprints. It is implemented for radio environment mining tasks.
@@
  • MLP [37, 38]: Many works have employed Multi-Layer Perceptron (MLP) to model complex mapping relationships in communication problems. We implement an MLP for the task of radio environment sensing and beam management.
@@
  • CNN [24]: A CNN-based predictor for FDD systems is proposed in [24], where the prediction of time-frequency CSI data is treated as a two-dimensional image processing task. It contains ten convolutional layers, where the convolution kernel size is $3\times 3$. We implement it for processing channel reconstruction tasks.
@@
  • Transformer [23]: A transformer-based parallel channel predictor is proposed in [23] for TDD systems, aiming to mitigate error propagation issues. We implement it using $3$ encoders and $2$ decoders for processing channel reconstruction tasks.
@@ V-A3 Network and Training Parameters
In the simulation, we adopt the Multi-Task Adapter module with $N_{a}=8$ for input and output feature alignment. For the configuration of the MoE-LoRA fine-tuning method, we choose the number of experts to be $8$, and we set $r=8$ for each LoRA matrix. For the output module, as mentioned above, for the specific task, we use either a three-layer MLP with 768-dimensional features or a three-layer CNN with $33x$ 3 kernels for the feature process. and we employ only one single-layer fully connected network to align the output dimensions. The smallest version [43] of GPT-2 with $F=768$ feature dimension is adopted, the first $N_{L}=6$ layers of which are deployed. Both the warm-up and cosine annealing scheduler are employed to train LLM4WM. The first $50$ epochs serve as the warm-up phase, where the learning rate increases linearly from the minimum value of $1\times{10}^{-5}$ to $1\times{10}^{-3}$. During the subsequent training phases, the learning rate is dynamically adjusted using the cosine annealing scheduler. Additional hyperparameters for model training are presented in Tab. III.
@@ V-A4 Performance Metric
To provide a comprehensive evaluation of the proposed scheme’s performance, we introduce an additional metric, SE, which reflects the overall performance of the communication system. SE is a crucial metric that indicates the system’s achievable rate, thereby capturing the effectiveness of the communication. It is calculated by Eq. (4), where $\bm{h}_{t,k}$ is the actual CSI and $\bm{w}_{t}$ is obtained as Eq. (7) with predicted $\bm{h}_{t,k}$. The communication SNR is defined as $1/{\sigma_{n}^{2}}$ and set as $10$ dB.
@@ V-B1 Overall Performance
TABLE IV: Performance of LLM4WM and other baselines: For mmWave link, the maximum SE is $9.32$ $\cdot$$(\text{s}\cdot\text{Hz})^{-1}$; For sub-6G link, the maximum SE is $6.33$ $\cdot$$(\text{s}\cdot\text{Hz})^{-1}$. The boldface denotes the highest score, while the underline marks the second-best result.
@@ V-B2 Generalization Experiments
TABLE V: Generalization Performance of LLM4WM and Other Baselines: the average loss across tasks is computed and presented in the final column labeled ”AVG.”. The boldface denotes the highest score, while the underline marks the second-best result.
@@
Generalization, referring to the ability of models to maintain performance in new communication scenarios, is crucial for real-world deployment as it reduces the need for frequent updates. We use only $10\%$ of the RMa dataset to transfer a model trained in the UMa scenario, as well as a model trained on the $1.9$ GHz sub-6 GHz link dataset to the $2.4$ GHz sub-6 GHz link dataset. Results in Tab. V indicate that, despite the challenges of multi-task generalization and transfer, our approach consistently outperforms others across most tasks. The slight dip in radio environment mining performance is due to the task’s relative simplicity in the LOS scenario, where smaller models like WiT excel. However, our model excels in more complex tasks like channel estimation, which requires understanding multidimensional features, further confirming that large models are better suited for dynamic real-world communication scenarios.
@@ V-B3 Hyper-parameter Analysis
To illustrate the rationale behind the hyperparameter settings, we conduct a thorough experiment on how hyperparameters affect the performance of LLM4WM. Specifically, we examine the effects of varying the LoRA rank and the number of experts, as depicted in Fig. 7. When we increase the LoRA rank while keeping the number of experts fixed at $8$, the performance of LLM4WM gradually improves. This can be attributed to the enhanced adaptability of the model to the data distribution due to the increase in trainable parameters. However, this improvement comes with higher training overhead as a cost. After weighing the trade-off between performance and computational efficiency, we determine that a LoRA rank of $8$ provided the optimal balance. Subsequently, with the LoRA rank fixed at its optimal value of $8$, we incrementally increase the number of experts. A similar trend to that of the LoRA rank was observed, as an increase in the number of experts effectively enhanced the model’s analytical and representational capacity. Balancing performance and computational efficiency, we determine that setting the number of experts to $8$ was the most appropriate choice.
@@ V-B4 Ablation Experiments
To assess the effectiveness of the proposed modules, ablation experiments are conducted by altering or removing the configurations of the multi-task adapter and backbone LLM modules. Variations for the multi-task adapter include w/o $\text{Adapter}_{\text{in}}$ (placing adapter only on the output side of the LLM), w/o $\text{Adapter}_{\text{out}}$ (placing adapter only on the input side of the LLM), and w/o Adapter (no adapters). For the backbone LLM, variations include w/o LLM (removing the large model) and frozen LLM (freezing pre-trained weights). Results in Table V show that all ablation configurations led to performance declines, highlighting the effectiveness of both the multi-task adapter and backbone LLM modules. Notably, the impact of removing the backbone LLM is significantly greater, indicating its critical role in the success of multi-task joint learning for wireless tasks.
@@ V-B5 Efficiency Evaluation
We evaluate the model’s training and inference cost of LLM4WM with other baselines to assess the difficulty of deploying the model in practical scenarios, as shown in Tab. VII. The same machine with $4$ Intel Xeon Platinum 8375C CPUs, $4$ NVIDIA GeForce RTX4090 GPUs, and $256$ GB of RAM is used to conduct all evaluations. To facilitate representation and comparison, the data in the table reflects the average performance across tasks for each method. Notably, the MoE-LoRA fine-tuning method results in LLM4WM having trainable parameters comparable to those of smaller models. This highlights both the training efficiency and parameter efficiency of LLM4WM, as adding new tasks increases the model’s parameters by only about $1.13$ M, a small fraction of the total. Additionally, utilizing a lightweight backbone LLM ensures that the overall inference speed of LLM4WM remains acceptable. Thus, LLM4WM demonstrates significant potential for deployment in future communication scenarios marked by increasing demand and the need for customized services involving numerous tasks.